# RP1M: A Large-Scale Motion Dataset for Piano Playing with Bi-Manual Dexterous Robot Hands

### v1

@@ Abstract
It has been a long-standing research goal to endow robot hands with human-level dexterity. Bi-manual robot piano playing constitutes a task that combines challenges from dynamic tasks, such as generating fast while precise motions, with slower but contact-rich manipulation problems. Although reinforcement learning based approaches have shown promising results in single-task performance, these methods struggle in a multi-song setting. Our work aims to close this gap and, thereby, enable imitation learning approaches for robot piano playing at scale. To this end, we introduce the Robot Piano 1 Million (RP1M) dataset, containing bi-manual robot piano playing motion data of more than one million trajectories. We formulate finger placements as an optimal transport problem, thus, enabling automatic annotation of vast amounts of unlabeled songs. Benchmarking existing imitation learning approaches shows that such approaches reach state-of-the-art robot piano playing performance by leveraging ${}^{~{}\diamond}$.
Keywords: Bi-manual dexterous robot hands, dataset for robot piano playing, imitation learning, robot learning at scale
@@ 1 Introduction
In this paper, we propose the Robot Piano 1 Million dataset (RP1M). This dataset comprises the motion data of high-quality bi-manual robot piano play. In particular, we train RL agents for each of the 2k songs and roll out each policy 500 times with different random seeds. To enable the generation of RP1M, we introduce a method to learn the fingering automatically by formulating finger placement as an optimal transport (OT) problem [8, 9]. Intuitively, the fingers are placed in a way such that the correct keys are pressed while the overall moving distance of the fingers is minimized. Agents trained using our automatic fingering match the performance of agents trained with human-annotated fingering labels. Besides, our method is easy to implement with almost no extra training time. The automatic fingering also allows learning piano playing with different embodiments, such as robots with four fingers only. With RP1M, it is now possible to train and test imitation learning approaches at scale. We benchmark various behavior cloning approaches and find that by using RP1M, existing methods reach state-of-the-art performance in terms of generalization capabilities in multi-song piano play. This work contributes in various ways:
@@
  • Using RP1M, we benchmark various approaches to robot piano playing, whereby existing imitation learning approaches reach new state-of-the-art results in motion synthesis on novel music pieces due to scaling with RP1M.
@@ 2 Related Work
Due to the difficulty of actually modeling the dynamics of the dexterous robot hand, recent methods resort to learning-based approaches, especially RL, which has achieved huge success in both robotics [19, 20, 1] and computer graphics [21]. To ease the training of dexterous robot hands with a large number of degrees of freedom (DoFs), demonstrations are commonly used [22, 23, 24, 25]. Due to the advance of both RL algorithms and simulation, recent work shows impressive results on dexterous hand manipulation tasks without human demonstrations. Furthermore, the policy trained in the simulator can further be deployed on real dexterous robot hands via sim-to-real transfer [26, 27, 28, 29, 30, 31].
@@
Piano Playing with Robots   Piano playing with robot hands has been investigated for decades. It is a challenging task since bi-manual robot hands should precisely press the right keys at the right time, especially considering its high-dimensional action space. Previous methods require specific robot designs [32, 33, 34, 35, 36, 37] or trajectory pre-programming [38, 39]. Recent methods enable piano playing with dexterous hands through planning [40] or RL [41] but are limited to simple music pieces. RoboPianist [4] introduces a benchmark for robot piano playing and demonstrates strong RL performance, but requires human fingering labels and performs worse in multi-task learning.
@@
Human fingering informs the agent of the correspondence between fingers and pressed keys at each time step. These labels require expert annotators and are, therefore, expensive to acquire in practice. Several approaches learn fingering from human-annotated data with different machine learning methods [6, 42, 43]. Moryossef et al. [44] extract fingering from videos to acquire fingering labels cheaply. Ramoneda et al. [45] propose to treat piano fingering as a sequential decision-making problem and use RL to calculate fingering but without considering the model of robot hands. Shi et al. [46] automatically acquires fingering via dynamic programming, but the solution is limited to simple tasks. In our paper, we do not introduce a separate fingering model, instead, similar to human pianists, fingering is discovered automatically while playing the piano.
@@
Datasets for Dexterous Robot Hands   Most large-scale datasets of dexterous robot hands focus on grasping various objects. To get suitable grasp positions, some methods utilize planners [47, 48, 49], while others use learned grasping policies [29], or track grasping motions of humans and imitate these motions on a robot hand [50]. Compared to the abundance of datasets for grasping, there exist relatively few datasets for object manipulation with dexterous robot hands. The D4RL benchmark [51] provides small sets of expert trajectories for four such tasks, consisting of human demonstrations and rollouts of trained policies. Zhao et al. [52] provide a small object manipulation dataset that utilizes a low-cost bi-manual platform with simple parallel grippers. Chen et al. [53] collect offline datasets for two simulated bi-manual manipulation tasks with dexterous hands. Table 1 summarizes the characteristics of these existing datasets. To the best of our knowledge, our RP1M dataset is the first large-scale dataset of dynamic, bi-manual manipulation with dexterous robot hands.
Table 1: Existing datasets on dexterous or bi-manual robotic manipulation.
@@
Dataset Task Dexterous hands Bi-manual Dynamic tasks Demonstra- tions
@@ 3 Background
Playing Piano with RL   We use RL to train specialist agents per song to control the bi-manual dexterous robot hands to play the piano. We frame the piano playing task as a finite Markov Decision Process (MDP). At time step $t$, the agent $\pi_{\theta}(a_{t}|s_{t})$, parameterized by $\theta$, receives state $s_{t}$ and takes action $a_{t}$ to interact with the environment and receives new state $s_{t+1}$ and reward $r_{t}$. The state and action spaces are described above and the reward $r_{t}$ gives an immediate evaluation of the agent’s behavior. We will introduce reward terms used for training in Section 4.1. The agent’s goal is to maximize the expected cumulative rewards over an episode of length $H$, defined as $\mathcal{J}=\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{t=0}^{H}\gamma^{t}r_{t}(s_{t},a_{t})\Big{]},$ where $\gamma$ is a discount factor ranging from 0 to 1.
@@ 4.1 Piano Playing without Human Fingering Labels
$K_{t}$ represents the set of keys to press at time step $t$ and $F$ represents the fingers of the robot hands. $c_{t}(k^{i},f^{j})$ represents the cost of moving finger $f^{j}$ to piano key $k^{i}$ at time step $t$ calculated by their Euclidean distance. $w_{t}(k^{i},f^{j})$ is a boolean weight. In our case, it enforces that each key in $K_{t}$ will be pressed by only one finger in $F$, and each finger presses at most one key. The constrained optimization problem in Eq. 1 is an optimal transport problem. Intuitively, it tries to find the best ”transport” strategy such that the overall cost of moving (a subset of) fingers $F$ to keys $K_{t}$ is minimized. We solve this optimization problem with a modified Jonker-Volgenant algorithm [55] from SciPy [56] and use the optimal combinations $(i^{*},j^{*})$ as the fingering for the agent.
We define a reward $r_{t}^{\text{OT}}$ based on $d_{t}^{\text{OT}}$ to encourage the agent to move the fingers close to the keys $K_{t}$, which is defined as:
@@
$c$ is a constant scale value and we use the same value as Tassa et al. [57]. The overall reward function is defined as:
@@
$r^{\text{Press}}$ and $r^{\text{Sustain}}_{t}$ represent the reward for correctly pressing the target keys and the sustain pedal. $r^{\text{Collision}}_{t}$ encourages the agent to avoid collision between forearms and $r^{\text{Energy}}_{t}$ prefers energy-saving behaviors. $\alpha_{1}$ and $\alpha_{2}$ are coefficient terms, and $\alpha_{1}=0.5$ and $\alpha_{2}=5\cdot 10^{-3}$ are adopted. Our method is compatible with any RL methods, and we use DroQ [58] in our paper. During RL training, the generated fingering will keep updating according to the state of hands. After giving an initial guess at the beginning, the OT-based fingering will be revised if the agent discovers a better fingering solution.
@@ 4.2 Analysis of Specialist RL Agents
Results   In Fig. 2, we compare our method with RoboPianist-RL both with and without human fingering. We use the same DroQ algorithm with the same hyperparameters for all experiments. RoboPianist-RL includes human fingering in its inputs, and the fingering information is also used in the reward function to force the agent to follow this fingering. Our method, marked as OT, removes the fingering from the observation space and uses OT-based finger placement to guide the agent to discover its own fingering. We also include a baseline, called No Fingering, that removes the fingering entirely. The first two columns of Fig. 2 show that our method without human-annotated fingering matches RoboPianst-RL’s performance on two different songs. Our method outperforms the baseline without human fingering by a large margin, showing that the proposed OT-based finger placement boosts the agent learning. The proposed method works well even on challenging songs. We test our method on Flight of the Bumblebee and achieve 0.79 F1 score after 3M training steps. To the best of our knowledge, we are the first to play the challenging song Flight of the Bumblebee with general-purpose bi-manual dexterous robot hands.
@@ 4.3 RP1M Dataset
To facilitate the research on dexterous robot hands, we collect and release a large-scale motion dataset for piano playing. Our dataset includes $\sim$1M expert trajectories covering $\sim$2k musical pieces. For each musical piece, we train an individual DroQ agent with the method introduced in Section 4.1 for 8 million environment steps and collect 500 expert trajectories with the trained agent. We chunk each sheet music every 550 time steps, corresponding to 27.5 seconds, so that each run has the same episode length. The sheet music used for training is from the PIG dataset [6] and a subset (1788 pieces) of the GiantMIDI-Piano dataset [7]. The stored episodes are converted to RLDS format [61].
@@ 5 Benchmarking Results
Baselines   We evaluated Behavior Cloning (BC) [62], Implicit Behavioral Cloning (IBC) [63], BC with a Recurrent Neural Network policy (BC-RNN) [64], and Diffusion Policy [5]. BC directly learns a policy by using supervised learning on observation-action pairs from expert demonstrations. IBC learns an implicit policy as an energy-based model conditioned on observation and action. BC-RNN uses an RNN as the policy network to encode a history of observations. Diffusion Policy learns to model the action distribution by inverting a process that gradually adds noise to a sampled action sequence. We used a CNN-based Diffusion Policy with DDIM [65] as the sampler. We use the same code and hyperparameters as Chi et al. [5].
Experiment Setup   We first train the policies with 3 different sizes of expert data: 50, 150, and 300 songs. We then evaluate the trained policies on 3 different groups of music pieces. (1) In-distribution songs: music pieces that overlap with the training sets. It shows the multitasking performance of the trained policies. (2) Easy out-of-distribution (OOD) songs: simple music pieces that do not overlap with the training songs. Those pieces are easy to play, with only slow motions and short horizons. (3) Hard out-of-distribution songs: difficult music pieces that do not overlap with the training songs. They contain more diverse motions and longer horizons. The out-of-distribution evaluation assesses the zero-shot generalization ability of the trained policies. We report the average F1 scores of each group of music pieces for policies trained with each baseline method.
@@
BC
Task #music 256 1024 4096 IBC BC-RNN Diffusion Policy
50 0.086 0.621 0.200 0.120 0.174 0.706
150 0.124 0.245 0.176 0.088 0.121 0.578
In-Dist. 300 0.084 0.249 0.102 0.083 0.025 0.596
50 0.06 0.278 0.109 0.128 0.204 0.446
150 0.085 0.257 0.172 0.058 0.050 0.465
Easy OOD 300 0.056 0.236 0.112 0.056 0.039 0.486
50 0.141 0.244 0.217 0.183 0.221 0.303
150 0.155 0.275 0.229 0.148 0.105 0.432
Hard OOD 300 0.145 0.253 0.181 0.144 0.081 0.440
Discussion As shown in Table 2, most baselines including BC (256), IBC, and BC-RNN have worse performance. This is because of the limited model capacity. We increase the model capacity of the BC baseline by increasing the hidden dimension of a 3-layer MLP from 256 to 1024 and we observe clear improvement in both in-distribution evaluation and OOD evaluation. However, when further increasing the model size to a 6-layer MLP with 4096 hidden dimensions, a performance drop is observed, meaning directly increasing the model capacity of MLP causes issues for model training. Among the baselines we have evaluated, Diffusion Policy performs the best in all cases, demonstrating that it is a strong baseline for the piano-playing task. Furthermore, when increasing the size of the training set, the OOD performance gradually increases. However, for in-distribution evaluation, Diffusion Policy still has a lower F1 score than our RL specialist and we observed a performance drop when increasing the dataset size, which indicates the Diffusion Policy has issues fitting our dataset well. This can be caused by multiple reasons, e.g., the limited model capacity or improper hyperparameters but we leave it as future work.

### v2

# RP1M: A Large-Scale Motion Dataset for Piano Playing with Bimanual Dexterous Robot Hands

@@ Abstract
It has been a long-standing research goal to endow robot hands with human-level dexterity. Bimanual robot piano playing constitutes a task that combines challenges from dynamic tasks, such as generating fast while precise motions, with slower but contact-rich manipulation problems. Although reinforcement learning-based approaches have shown promising results in single-task performance, these methods struggle in a multi-song setting. Our work aims to close this gap and, thereby, enable imitation learning approaches for robot piano playing at scale. To this end, we introduce the Robot Piano 1 Million (RP1M) dataset, containing bimanual robot piano playing motion data of more than one million trajectories. We formulate finger placements as an optimal transport problem, thus, enabling automatic annotation of vast amounts of unlabeled songs. Benchmarking existing imitation learning approaches shows that such approaches reach promising robot piano playing performance by leveraging ${}^{~{}\diamond}$.
Keywords: Bimanual dexterous robot hands, dataset for robot piano playing, imitation learning, robot learning at scale
@@ 1 Introduction
In this paper, we propose the Robot Piano 1 Million dataset (RP1M). This dataset comprises the motion data of high-quality bi-manual robot piano play. In particular, we train RL agents for each of the 2k songs and roll out each policy 500 times with different random seeds. To enable the generation of RP1M, we introduce a method to learn the fingering automatically by formulating finger placement as an optimal transport (OT) problem [8, 9]. Intuitively, the fingers are placed in a way such that the correct keys are pressed while the overall moving distance of the fingers is minimized. Agents trained using our automatic fingering match the performance of agents trained with human-annotated fingering labels. Besides, our method is easy to implement with almost no extra training time. The automatic fingering also allows learning piano playing with different embodiments, such as robots with four fingers only. With RP1M, it is now possible to train and test imitation learning approaches at scale. We benchmark various behavior cloning approaches and find that using RP1M, existing methods perform better in terms of generalization capabilities in multi-song piano play. This work contributes in various ways:
@@
  • Using RP1M, we benchmark various approaches to robot piano playing, whereby existing imitation learning approaches reach promising results in motion synthesis on novel music pieces due to scaling with RP1M.
@@ 2 Related Work
Piano Playing with Robots   Piano playing with robot hands has been investigated for decades. It is a challenging task since bimanual robot hands should precisely press the right keys at the right time, especially considering its high-dimensional action space. Previous methods require specific robot designs [10, 11, 12, 13, 14, 15] or trajectory pre-programming [16, 17]. Recent methods enable piano playing with dexterous hands through planning [18] or RL [19] but are limited to simple music pieces. RoboPianist [4] introduces a benchmark for robot piano playing and demonstrates strong RL performance, but requires human fingering labels and performs worse in multi-task learning.  Yang et al. [20] improves the policy training performance by considering the bionic constraints of the anthropomorphic robot hands.
@@
Human fingering informs the agent of the correspondence between fingers and pressed keys at each time step. These labels require expert annotators and are, therefore, expensive to acquire in practice. Several approaches learn fingering from human-annotated data with different machine learning methods [6, 21, 22]. Moryossef et al. [23] extract fingering from videos to acquire fingering labels cheaply. Ramoneda et al. [24] proposes to treat piano fingering as a sequential decision-making problem and use RL to calculate fingering but without considering the model of robot hands. Shi et al. [25] automatically acquires fingering via dynamic programming, but the solution is limited to simple tasks. Concurrent work [26] obtains fingering labels from YouTube videos and trains a diffusion policy to play hundreds of songs. In our paper, we do not introduce a separate fingering model, instead, similar to human pianists, fingering is discovered automatically while playing the piano, hereby largely expanding the pool of usable data to train a generalist piano-playing agent.
@@
Datasets for Dexterous Robot Hands   Most large-scale datasets of dexterous robot hands focus on grasping various objects. To get suitable grasp positions, some methods utilize planners [27, 28, 29], while others use learned grasping policies [30], or track grasping motions of humans and imitate these motions on a robot hand [31]. Compared to the abundance of datasets for grasping, there exist relatively few datasets for object manipulation with dexterous robot hands. The D4RL benchmark [32] provides small sets of expert trajectories for four such tasks, consisting of human demonstrations and rollouts of trained policies. Zhao et al. [33] provides a small object manipulation dataset that utilizes a low-cost bimanual platform with simple parallel grippers. Chen et al. [34] collects offline datasets for two simulated bimanual manipulation tasks with dexterous hands. Furthermore, Fan et al. [35] proposes a large-scale dataset for bimanual hand-object manipulation with human hands rather than robot hands. Table 1 summarizes the characteristics of these existing datasets. To the best of our knowledge, our RP1M dataset is the first large-scale dataset of dynamic, bimanual piano playing with dexterous robot hands.
We further discuss related work on dexterous robot hands and generalist agents in Appendix A.
Table 1: Existing datasets on dexterous or bimanual robotic manipulation.
@@
Dataset Task Dexterous hands Bimanual Dynamic tasks Demonstra- tions
@@ 3 Background
Playing Piano with RL   We use RL to train specialist agents per song to control the bimanual dexterous robot hands to play the piano. We frame the piano playing task as a finite Markov Decision Process (MDP). At time step $t$, the agent $\pi_{\theta}(a_{t}|s_{t})$, parameterized by $\theta$, receives state $s_{t}$ and takes action $a_{t}$ to interact with the environment and receives new state $s_{t+1}$ and reward $r_{t}$. The state and action spaces are described above and the reward $r_{t}$ gives an immediate evaluation of the agent’s behavior. We will introduce reward terms used for training in Section 4.1. The agent’s goal is to maximize the expected cumulative rewards over an episode of length $H$, defined as $\mathcal{J}=\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{t=0}^{H}\gamma^{t}r_{t}(s_{t},a_{t})\Big{]},$ where $\gamma$ is a discount factor ranging from 0 to 1.
@@ 4.1 Piano Playing without Human Fingering Labels
$K_{t}$ represents the set of keys to press at time step $t$ and $F$ represents the fingers of the robot hands. $c_{t}(k^{i},f^{j})$ represents the cost of moving finger $f^{j}$ to piano key $k^{i}$ at time step $t$ calculated by their Euclidean distance. $w_{t}(k^{i},f^{j})$ is a boolean weight. In our case, it enforces that each key in $K_{t}$ will be pressed by only one finger in $F$, and each finger presses at most one key. The constrained optimization problem in Eq. 1 is an optimal transport problem. Intuitively, it tries to find the best ”transport” strategy such that the overall cost of moving (a subset of) fingers $F$ to keys $K_{t}$ is minimized. We solve this optimization problem with a modified Jonker-Volgenant algorithm [37] from SciPy [38] and use the optimal combinations $(i^{*},j^{*})$ as the fingering for the agent. The fingering is calculated on the fly based on the hands’ state, so during the RL training, the fingering adjusts according to the robot hands’ state.
We define a reward $r_{t}^{\text{OT}}$ to encourage the agent to move the fingers close to the keys $K_{t}$. which is defined as:
@@
$c$ is a constant scale value as used in Tassa et al. [39] and $d_{t}^{\text{OT}}$ is the distance between fingers and target keys obtained by solving Eq. 1. $r_{t}^{\text{OT}}$ increases exponentially as $d_{t}^{\text{OT}}$ decreases and is set as 1 once $d_{t}^{\text{OT}}$ is smaller than a pre-defined threshold (0.01). The overall reward function is defined as:
@@
$r^{\text{Press}}$ and $r^{\text{Sustain}}_{t}$ represent the reward for correctly pressing the target keys and the sustain pedal. $r^{\text{Collision}}_{t}$ encourages the agent to avoid collision between forearms and $r^{\text{Energy}}_{t}$ prefers energy-saving behaviors. $\alpha_{1}$ and $\alpha_{2}$ are coefficient terms, and $\alpha_{1}=0.5$ and $\alpha_{2}=5\cdot 10^{-3}$ are adopted. Our method is compatible with any RL methods, and we use DroQ [40] in our paper.
@@ 4.2 Analysis of Specialist RL Agents
Results   In Fig. 2, we compare our method with RoboPianist-RL both with and without human fingering. We use the same DroQ algorithm with the same hyperparameters for all experiments. RoboPianist-RL includes human fingering in its inputs, and the fingering information is also used in the reward function to force the agent to follow this fingering. Our method, marked as OT, removes the fingering from the observation space and uses OT-based finger placement to guide the agent to discover its own fingering. We also include a baseline, called No Fingering, that removes the fingering entirely. The first two columns of Fig. 2 show that our method without human-annotated fingering matches RoboPianst-RL’s performance on two different songs. Our method outperforms the baseline without human fingering by a large margin, showing that the proposed OT-based finger placement boosts the agent learning. The proposed method works well even on challenging songs. We test our method on Flight of the Bumblebee and achieve 0.79 F1 score after 3M training steps. To the best of our knowledge, we are the first to play the challenging song Flight of the Bumblebee with general-purpose bimanual dexterous robot hands.
@@ 4.3 RP1M Dataset
To facilitate the research on dexterous robot hands, we collect and release a large-scale motion dataset for piano playing. Our dataset includes $\sim$1M expert trajectories covering $\sim$2k musical pieces. For each musical piece, we train an individual DroQ agent with the method introduced in Section 4.1 for 8 million environment steps and collect 500 expert trajectories with the trained agent. We chunk each sheet music every 550 time steps, corresponding to 27.5 seconds, so that each run has the same episode length. The sheet music used for training is from the PIG dataset [6] and a subset (1788 pieces) of the GiantMIDI-Piano dataset [7].
@@ 5 Benchmarking Results
Baselines   We evaluated Behavior Cloning (BC) [43], Behavior Transformer (BeT) [44], Diffusion Policy [5] with U-Net (DP-U) [45] and with Transformer (DP-T) [46]. BC directly learns a policy by using supervised learning on observation-action pairs from expert demonstrations. BeT clusters continuous actions into discrete bins using k-means, allowing it to model high-dimensional, continuous, multimodal action distributions as categorical distributions [44]. Diffusion Policy learns to model the action distribution by inverting a process that gradually adds noise to a sampled action sequence. We evaluated both the CNN-based (U-Net) Diffusion Policy (DP-U) and the Transformer-based Diffusion Policy (DP-T) with DDPM [47]. We use the same code and hyperparameters as Chi et al. [5]. Detailed descriptions of the baselines as well as hyperparameters are given in Section C.1.
Experiment Setup   We train the policies on subsets of the RP1M dataset with different sizes: 12, 25, 50, 100, 150. We then evaluate the trained policies on both i) 12 in-distribution songs: music pieces that overlap with the training sets, and ii) 20 out-of-distribution (OOD) songs: music pieces that do not overlap with the training songs. The selected songs are very challenging and contain diverse motions and long horizons. In the experiment, we report zero-shot evaluation results without fine-tuning. We report the average F1 scores of each group of music pieces for policies trained with each baseline method. We list the selected songs for evaluation in Section C.2.
@@
In-disribution Out-of-distribution
# music 12 25 50 100 150 12 25 50 100 150
BC-MLP 0.529 0.315 0.319 0.193 0.250 0.079 0.119 0.100 0.108 0.187
BeT 0.062 0.080 0.065 0.078 0.088 0.094 0.110 0.111 0.120 0.125
DP-U 0.539 0.541 0.546 0.505 0.454 0.181 0.189 0.198 0.215 0.256
DP-T 0.357 0.304 0.297 0.301 0.318 0.186 0.210 0.230 0.291 0.316
Discussion We present the benchmarking performance of multi-task agents in Table 2. For the in-distribution evaluation, compared to F1 scores obtained with our RL specialist agents in Fig. 4, we notice a performance gap across all baselines. This gap widens as the data size increases. When trained on a smaller dataset with 12 training songs, DP-U performs comparably to BC-MLP and slightly outperforms DP-T, while BeT experiences a significant performance drop. This decline may be attributed to hyperparameter choices, such as the number of action bins. Although we used the same number of action bins as the official implementation, the complexity of our tasks suggests that this configuration may be inadequate, and increasing the number of bins could improve performance.
As the dataset size increases, we observe that Diffusion Policy outperforms the other baselines. DP-U and DP-T show performance drops of 15.77% and 10.92%, respectively, while BC-MLP suffers a more significant decline of 52.74%. Similar performance gaps have been noted in previous work [4] and concurrent research [26] suggests a hierarchical policy structure, although it still lags behind RL specialists. This highlights the need for future research to address the performance gap between RL specialists and multi-task agents.
In the zero-shot out-of-distribution evaluation, we find that performance improves for all evaluated baselines as the training data size increases. Specifically, the F1 scores for DP-U and DP-T rise from 0.181 to 0.256 and from 0.186 to 0.316, respectively, when the number of training songs is increased from 12 to 150. This suggests that larger datasets enhance the generalization capabilities of multi-task agents. We hope that releasing our large-scale RP1M dataset will contribute to the development of robust generalist piano-playing agents within the research community.